OpenStudy (anonymous):

SOMEONE HELP 8th GRADE PRE-ALGEBRA

OpenStudy (snuggielad):

Sam, lets not condone cheating.

OpenStudy (anonymous):

if u want to learn or cheat, mathway.com

OpenStudy (anonymous):

lol

OpenStudy (snuggielad):

Class, this is one point where I can tell you cheating is pointless even if it is easier. What is your question?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

m+1 1/3=5

OpenStudy (snuggielad):

Are you missing a sign?

OpenStudy (snuggielad):

Double check and be sure buddy it seems like you are. That or you subtract 1 1/3 from both sides. Use the Equation tab below below the reply box.

OpenStudy (anonymous):

no its 1 and 1 3rd

OpenStudy (anonymous):

\[m+1 1/3=5\]

OpenStudy (calculusxy):

You want to find the value of m?

OpenStudy (anonymous):

@calculusxy yes

OpenStudy (snuggielad):

Subtract 1 and 1 third from both sides.

OpenStudy (anonymous):

its about solving systems of equations

OpenStudy (snuggielad):

5-1 1/3=?

OpenStudy (anonymous):

8th grade math

OpenStudy (snuggielad):

\[5-1\frac{ 1 }{ 3 }=3.67\]

OpenStudy (calculusxy):

@classof2018 Is there an option of -9.34?

OpenStudy (snuggielad):

The goal is to isolate the variable on one side. When you move one number over the equal sign you ALWAYS change the sign. For example if you move a +3 to the other side of the equal sign you make it a -3 which means subtract and if you move a -3 to the other side of the equal sign you make it a +3

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

lol classof2018 nty lmfao

OpenStudy (snuggielad):

Calc, It is m + 1 and 1/3=5

OpenStudy (snuggielad):

or better written \[m+1\frac{ 1 }{ 3 }=5\]

OpenStudy (calculusxy):

i would make 1 1/3 into decimal 1 1/3=1.33 Then it would be m+1.33=5 FInaly, 5-1.33=3.67 m=3.67